OpenStudy (nurali):

A list of numbers has an average of 48 and an SD of 12. The list is transformed by subtracting 3 from each entry and then dividing the result by 5.

OpenStudy (nurali):

Find the average of the transformed list.

OpenStudy (nurali):

@phi

OpenStudy (phi):

the average is sum/N we don't know the sum or N but we know sum/N= 48 and sum= N*48 if we subtract 3 from each number, and there are N numbers, how much do we subtract? I think N*3 so the new sum is Sum_new= N*48-N*3= N(48-3)= N*45 if we divide Sum_new by N we get the new average

OpenStudy (phi):

if we toss in divide by 5 then Sum_new= N*45/5 = N*9 Sum_new/N= 9 is the average for this new list

OpenStudy (nurali):

thanks

OpenStudy (nurali):

please help @phi

OpenStudy (phi):

how do you find the s.d. of a list?

OpenStudy (nurali):

A list of numbers has an average of 48 and an SD of 12. The list is transformed by subtracting 3 from each entry and then dividing the result by 5.

OpenStudy (nurali):

Find the SD of the transformed list.

OpenStudy (phi):

yes. How do you find the standard deviation of a list ?

OpenStudy (nurali):

please help.

OpenStudy (phi):

to find the standard deviation squared (the variance) of N numbers: \[\sigma^2=\frac{ 1 }{ N }\sum_{i=1}^{N}(x_i-\bar{x})^2\]

OpenStudy (phi):

the original list has s.d.=12 and var= 12^2= 144 We don't know how many numbers we have, but the computation would look like this: \[ \frac{(a-48)^2}{N} + \frac{(b-48)^2}{N}+...+\frac{(z-48)^2}{N}= 144 \] we can write this as \[ (a-48)^2 + (b-48)^2 + ... + (z-48)^2 = 144 N \]

OpenStudy (phi):

The transformed numbers will each have the form \[ \frac{a-3}{5} \] Their variance (that is, standard deviation squared) is computed as \[ \frac{1}{N}\left(\frac{a-3}{5}-9\right)^2 + \frac{1}{N}\left(\frac{b-3}{5}-9\right)^2 + ... + \frac{1}{N}\left(\frac{z-3}{5}-9\right)^2= \sigma^2\] notice that we use the mean of the transformed list (the 9) rather than the mean of the original list. we can simplify this a bit. Multiply both sides by N. Write 9 as 45/5 so we can combine the fractions: \[ \left(\frac{a-3-45}{5}\right)^2 + \left(\frac{b-3-45}{5}\right)^2+...+ \left(\frac{z-3-45}{5}\right)^2= N\sigma^2\] we now get \[ \frac{(a-48)^2}{25}+ \frac{(b-48)^2}{25}+ ... +\frac{(z-48)^2}{25}= N\sigma^2\] or, after multiplying both sides by 25: \[ (a-48)^2 + (b-48)^2 + ...+(c-48)^2= 25N\sigma^2 \] Interestingly, we know the sum on the left side, it is 144N, so we have \[ 144N= 25N\sigma^2 \] solving for sigma, we get \[ \sigma= 12/5 \]

OpenStudy (nurali):

Thanks